OpenStudy (anonymous):

Let f(x)=2x^-1, g(x)=2x+5 and H(x)=x-3. Perform the indicated operation. g(h(x))

OpenStudy (amistre64):

plug in h(x) in the x spot of g like this: g(h(x)) = 2(h(x)) +5 = 2(x-3) +5 = 2x -6 +5 g(h(x)) = 2x -1

OpenStudy (anonymous):

so 21 is the answer

OpenStudy (anonymous):

2x-1 sorry

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

thanks